| 12/23/2004 | ||||||||||
| 0:09 | Cal Mb5 Hv ramping upto 1000 | TG | ||||||||
| 0:15 | Cal Mb5 Hv ramping upto 1250 | TG | ||||||||
| 0:25 | Cal Mb5 Hv ramping upto 1250 | TG | ||||||||
| 0:35 | Cal Mb5 Hv ramping upto 1500 | TG | ||||||||
| 0:52 | Cal Mb5 Hv ramping upto 1750 | TG | ||||||||
| 1:00 | Cal Mb5 Hv ramping upto 2000 | TG | ||||||||
| 1:43 | Cal trigger rate 75 | TG | ||||||||
| 2:20 | Cal trigger rate 70 | TG | ||||||||
| 7:27 | manual restart cdps2 and cdps3 | HP | ||||||||
| 7:34 | Receiving intermittance data due to satellite | KC | ||||||||
| Satellite is screened by horizon | ||||||||||
| We will receive data in about 20 minutes | ||||||||||
| 7:55 | Receiving good data | KC | ||||||||
| 8:09 | ESC tell us we are receiving mim criteria data. | KC | ||||||||
| 12:23 | standby mode by sonny | JH | ||||||||
| 12:25 | scd_ch_coher.stol | |||||||||
| scd_ch_mask_dummy.stol | ||||||||||
| scd_ch_mask_hot2.stol | ||||||||||
| scd_ch_mask_hot3.stol | ||||||||||
| scd_ch_mask_hot.stol | ||||||||||
| daq_restart.stol | ||||||||||
| BUT the commanding window closed while uploading scd_ch_mask_hot2.stol. | ||||||||||
| 12:56 | TCD setting for item 3 and 7changed from 300 to 200 according to Jim's request via email | YS | ||||||||
| daq_restart.stol | ||||||||||
| 14:19 | changed SCD #2880 threshold from 40000 to 65000 | JH | ||||||||
| 15:03 | changed all of SCD hot channel value to 65535 | JH | ||||||||
| 16:43 | Cal trigger rate 68 | KC | ||||||||
| 16:56 | Data disconnected due to switching satellite | KC | ||||||||
| 17:10 | Data resumed | KC | ||||||||
| 17:32 | ESC started uploading pmf013.* | |||||||||
| 18:30 | CDPS3 shut down suddenly | |||||||||
| 18:36 | ESC completed uploading | JH | ||||||||
| SCI_SW_LOGFILE_SEND | ||||||||||
| SCI_SW_DISK_LIST | ||||||||||
| 18:39 | SCI_SW_FILE_LIST | |||||||||
| SCI_SW_CUSTOM | ||||||||||
| 19:23 | getting play-back data | JH | ||||||||
| 20:10 | Z-HI trigger rate was too low, also Network rate was between 1~2 | JH | ||||||||
| TCD power restart | ||||||||||
| 20:58 | DAQ power restart | |||||||||
| commands went through well, we checked with ESC | ||||||||||
| but it doesn't work. | ||||||||||
| 22:05 | Power cycling of TCD and Resetting DAQ according to sonny's manual | JH | ||||||||
| TCD trigger rate was recovered but there was a new problem on CAL, whose trigger rate was about 20, too high .. | SO | |||||||||
| There were some changes on the manual, especailly chapter 3. Power cycling of TCD, chapter 4 Resetting DAQ, chapter 5 Rebooting Science Flight computer. Please refer to the manual | SO | |||||||||
| 22:15 | As Sonny`s recommend, we sent scripts on Chapter 4.1, 4.5, 4.6 in order, but it didn`t work properly. | SO | ||||||||
| CAL trigger threshold 68 | ||||||||||
| 22:23 | We altered the Cal`s thresold to 75. As a result, its rate was about 0.7 | SO | ||||||||